OpenStudy (anonymous):

Volume= 18 cm3 Height of the pyramid= 3 cm B1= 2 cm B2= 3 cm What is the height of the base?

OpenStudy (anonymous):

you mean area of the base?

OpenStudy (anonymous):

wait, no. base area is .5xh(b1+b2) so i need the height

OpenStudy (anonymous):

i though height was given as 3 cm?

OpenStudy (anonymous):

thats the height of the pyramid

OpenStudy (anonymous):

okay, and what do you need to evaluate?

OpenStudy (anonymous):

the height of the base

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Pyramid_(geometry) are you talking about that? how is there a height of the base?

OpenStudy (anonymous):

x=18cm3

OpenStudy (anonymous):

ah, so you should say its a trapezoidal pyramid.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay, do you know the formula to the volume of a pyramid?

OpenStudy (anonymous):

yes. 1/3 x B x H

OpenStudy (anonymous):

what is B in the formula?

OpenStudy (anonymous):

i dont know. b1 = 2 cm b2= 3 cm

OpenStudy (anonymous):

yess.

OpenStudy (anonymous):

okay you must always know what the letters mean in a formula. for example the area of a rectangle is the its length multiplied by its width. you shouldn't remember it as l x b, because when I ask you what l and b are, you don't have an answer to that.

OpenStudy (anonymous):

volume of a pyramid is one-thirds the product of the base area and its height

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

so B is the area of the base. right, so substitute H and V in the equation

OpenStudy (anonymous):

right

OpenStudy (anonymous):

soooo,,,,,,,,,,,,,,,,,,,,

OpenStudy (anonymous):

so, tell me what you got after substituting the values of V and H.

OpenStudy (anonymous):

V= 18cm H= 3cm. h= i dont know

OpenStudy (anonymous):

substitute V = 18, H =3 in V = 1/3 x B x H and find the value of B.

OpenStudy (anonymous):

but H is the height of the pyramid, i need the height of the base (h)

OpenStudy (anonymous):

Yes, volume of a pyramid is 1/3 x Base area x Height of the pyramid. so once you find the base area of the pyramid, substitute its value as 1/2(b1+b2) x h and find h

OpenStudy (anonymous):

so how do you find h?

OpenStudy (anonymous):

I just told you. can you please just for the love of god substitute V and H in V = 1/3 B x H and give me an answer for B?

OpenStudy (anonymous):

BUT I DONT KNOW THE BASE. sorry i'm not good at math, my lord.

OpenStudy (anonymous):

what is so hard about this? 18 = 1/3 x B x 3 B =18 but we know that B = 1/2(b1+b2) x h so 18 = 1/2(2+3) x h h = 36/5

OpenStudy (anonymous):

you just have to work backward to get the base. you won't always find equations in the correct order. you must know how to manipulate equations to get your variable.

OpenStudy (anonymous):

where did we get 36

OpenStudy (anonymous):

multiply 18 by 2 and divide by 5.

OpenStudy (anonymous):

i got h = 7.2. is this correct?

OpenStudy (anonymous):

yeah. oooo you used a calculator to find the value of 36/5! good on you!

OpenStudy (anonymous):

btw, which grade are you in?

OpenStudy (anonymous):

7th

OpenStudy (anonymous):

okay. Have you ever done manipulation of equations?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

like if x = y+2, what is y?

OpenStudy (anonymous):

x-2?

OpenStudy (anonymous):

then why do you say you haven't done any algebraic manipulation before?

OpenStudy (anonymous):

btw, h= 2.4. 1/2 x 2.4 (3+2) = 6. 6x3=18.

OpenStudy (anonymous):

why do you say 6 x 3 = 18?

OpenStudy (anonymous):

is the formula for volume of a pyramid area of base x height?

OpenStudy (anonymous):

because the height of the pyramid is 3 cm, so to find the volume you have to multiply the base by the height.

OpenStudy (anonymous):

oh yeah, divide by 3. nevermind...

OpenStudy (anonymous):

good, at least you figured that out on your own.

OpenStudy (anonymous):

oh stopp

OpenStudy (anonymous):

:D